\begin{thebibliography}{99}

\bibitem{Sha} C.E.Shannon, A mathematical theory of communication, Bell Syst.Tech.J.,Vol.27(1948),
pp.379-423 and pp.623-656.
\bibitem{AD} J.Acz\'el and Z.Dar\'oczy, On measures of information and their characterizations, Academic Press, 1975.
\bibitem{Ren} A.R\'enyi, On measures of entropy and information, in Proc. 4th
Berkeley Symp., Mathematical and Statistical Probability, Berkeley,
CA: Univ. Calif. Press, Vol. 1(1961), pp. 547-561.
\bibitem{Tsa} C. Tsallis, Possible generalization of Bolzmann-Gibbs statistics, 
J.Stat. Phys., Vol. 52(1988), pp. 479-487.
\bibitem{Suy} H.Suyari, Generalization of Shannon-Khinchin axioms to nonextensive systems and the uniqueness theorem for the nonextensive entropy, 
IEEE Trans. Information Theory, Vol.50(2004), pp.1783-1787.
\bibitem{Csi} I. Csisz\'ar, Axiomatic characterizations of information measures, Entropy,Vol.10(2008),
pp.261-273.
%\bibitem{ST} B.D.Sharma and L.J.Taneja,Entropy of type $(\alpha,\beta)$ and other generalized measures in information theory,  Metrika, Vol.22(1975),pp.205-215.
%\bibitem{Mit} D.P.Mittal, On some functional equations concerning entropy, directed divergence and inaccuracy,Metrika, Vol.22(1975),pp.35-45.
\bibitem{BR} E.P.Borges and I.Roditi,A family of nonextensive entropies,
Phys.Lett.A, Vol.246(1998),pp.399-402. 
\bibitem{KLS1} G.Kaniadakis, M.Lissia and A.M.Scarfone,
Deformed logarithms and entropies,
 Physica A, Vol.340(2004),pp.41-49.
\bibitem{KLS2} G.Kaniadakis, M.Lissia and A.M.Scarfone,
Two-parameter deformations of logarithm, exponential, and entropy: A consistent framework for generalized statistical mechanics,
 Phys.Rev.E, Vol.71(2005),046128.
\bibitem{WS} T.Wada and H.Suyari,
A two-parameter generalization of Shannon-Khinchin axioms and the uniqueness theorem,
Phys.Lett.A, Vol.368(2007),pp.199-205.
\bibitem{Hob} A.Hobson, A new theorem of information theory, J.Stat.Phys.,Vol.1(1969),pp.383-391.
\bibitem{Furu1} S.Furuichi, On uniqueness theorems for Tsallis entropy and Tsallis relative entropy, 
IEEE Trans. Information Theory, Vol.51(2005), pp.3638-3645.  
\bibitem{FYK} S.Furuichi, K.Yanagi and K.Kuriyama, Fundamental properties of Tsallis relative entropy, J.Math.Phys.,Vol.45(2004),pp.4868-4877.
\bibitem{Furu2} S.Furuichi, Information theoretical properties of Tsallis entropies, J.Math.Phys., Vol.47(2006), 023302.
\bibitem{Furu3} S.Furuichi, On the maximum entropy principle and the minimization of the Fisher information in Tsallis statistics, J.Math.Phys., Vol.50(2009), 013303.
\bibitem{Csi2} I. Csisz\'ar, Information measures: A critical survey, Transactions of the Seventh Prague Conference on
Information Theory, Statistical Decision Functions, Random Processes 
and of the 1974 European meeting of Statisticians, 
held at Prague, from August 18 to 23, 1974 Volume B, pp.73-86, ACADEMIA, Publising House of the Czechoslovak Academy of Sciences, Prague, 1978.
(D. Reidel Publishing Company, DORDRECHT:HOLLAND/BOSTON:U.S.A.)
%\bibitem{AN} S.Amari and H.Nagaoka, Methods of information geometry, AMS Oxford University Press, 2000.

\bibitem{Ama} S.Amari, $\alpha$-Divergence Is Unique, Belonging to Both $f$-Divergence and Bregman Divergence Classes, IEEE Trans. Information theory, Vol. 55(2009), pp.4925-4931.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{graphicx}
\usepackage{amsfonts}
\usepackage{theorem}
\newtheorem{Lem}{Lemma}[section]
\newtheorem{Def}[Lem]{Definition}
\newtheorem{The}[Lem]{Theorem}
\newtheorem{Prop}[Lem]{Proposition}

\newtheorem{Ax}[Lem]{Axiom}
\newcommand{\qed}{\hbox{\rule{6pt}{6pt}}}

\begin{document}

% paper title
\title{An axiomatic characterization of a two-parameter extended relative entropy}

\author{Shigeru Furuichi$^1$\footnote{E-mail:furuichi@chs.nihon-u.ac.jp}\\
$^1${\small Department of Computer Science and System Analysis,}\\
{\small College of Humanities and Sciences, Nihon University,}\\
{\small 3-25-40, Sakurajyousui, Setagaya-ku, Tokyo, 156-8550, Japan}}
\date{}
\maketitle
{\bf Abstract.} The uniqueness theorem for a two-parameter extended relative entropy is proven.
This result extends our previous one, the uniqueness theorem for a one-parameter extended relative entropy,
to a two-parameter case.
In addition, the properties of a two-parameter extended relative entropy are studied.
\vspace{3mm}

{\bf Keywords : } Tsallis relative entropy, two-parameter extended entropy, axiomatic characterization and uniqueness theorem
\vspace{3mm}

{\bf 2000 Mathematics Subject Classification : }  94A17, 62B10 and 46N55 

\vspace{3mm}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%  Section1  %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}
Shannon entropy \cite{Sha} is one of fundamental quantities in classical information theory 
and uniquely determined by the 
Shannon-Khinchin axiom or the Faddeev axiom. 
One-parameter extensions for Shannon entropy have been studied by many researchers \cite{AD}.
The R\'enyi entropy \cite{Ren} and the Tsallis entropy \cite{Tsa} are famous.
In the paper \cite{Suy}, the uniqueness theorem for the Tsallis entropy was proven.
See also the paper \cite{Csi} and the references therein, for the axiomatic characterizations of one-parameter extended entropies.
The two-parameter family of entropy was introduced by Borges and Roditi in \cite{BR} by the use of
the generalized Jackson derivative method.
Recently, a two-parameter extended entropy, which is essentially same form with the two-parameter family of entropy in \cite{BR},  has studied by 
several researchers \cite{KLS1,KLS2,WS} and 
the uniqueness theorem for a two-parameter extended entropy was proven in \cite{WS} by generalizing
the Shannon-Khinchin axiom. 
In this paper, we denote a two-parameter extended entropy by
$$
S_{\alpha,\beta}(x_1,x_2,\cdots,x_n)= \sum_{j=1}^n \frac{x_j^{\alpha}-x_j^{\beta}}{\beta-\alpha},
\quad (\alpha \neq \beta)
$$
for two real numbers $\alpha$ and $\beta$ such that 
$0\leq \alpha \leq 1 \leq \beta$ or $0 \leq \beta \leq 1 \leq \alpha$.
If we take $\alpha = 1$ or $\beta =1$, then it recovers
the Tsallis entropy defined by 
$$
S_{q}(x_1,x_2,\cdots,x_n)\equiv \sum_{j=1}^n\frac{x_j-x_j^q}{q-1},\quad (1\neq q \geq 0).
$$
The Tsallis entropy recovers Shannon entropy
$$
S_1(X)\equiv -\sum_{j=1}^n x_j \log x_j 
$$
 in the limit $q \to 1$.

In this paper, we study on information measure (entropy) defined for two probability distributions.
The relative entropy (Kullback-Leibler information or divergence) is defined for {\it two} probability distributions
 $X=\left\{x_1,\cdots,x_n\right\}$ and $Y=\left\{y_1,\cdots,y_n\right\}$:
$$D_1(X\vert \vert Y)\equiv \sum_{j=1}^n x_j \left(\log x_j -\log y_j\right).$$
Since Shannon entropy is defined for {\it one} probability distribution
and it can be reproduced by the relative entropy as
$\log n -D_1(X\vert \vert U) $ for the uniform distribution $U=\left\{1/n,\cdots,1/n\right\}$,
 the relative entropy can be regarded as a generalization for Shannon entropy.
We here note that we have one-parameter extended relative entropies 
such as the R\'enyi relative entropy $D^{R}_q(X\vert Y)$, 
$\alpha$-divergence $D^{(\alpha)}(X\vert \vert Y)$ and the Tsallis entropy $D^{T}_q(X\vert \vert Y)$.
These are defined by
\begin{eqnarray*}
&&D^{R}_q(X \vert \vert  Y) \equiv \frac{1}{q-1}\log \sum_{j=1}^n x_j^q y_j^{1-q}, \\
&&D^{(\alpha)}(X\vert \vert Y) \equiv \frac{4}{1-\alpha^2}\left(1-\sum_{j=1}^n x_j^{\frac{1-\alpha}{2}}y_j^{\frac{1+\alpha}{2}}\right), \\
&&D^{T}_q(X\vert \vert Y) \equiv\sum_{j=1}^n  \frac{x_j- x_j^qy_j^{1-q}}{1-q}, 
\end{eqnarray*}
for $q \neq 1$ and $\alpha \neq \pm 1$.
These quantities recover the relative entropy in their limit $q \to 1$ or $\alpha \to \pm 1$.
These quantities are also essentially same one in the sense that
\begin{eqnarray*}
&&D^{(q)}(X\vert \vert Y) = \frac{1}{q}D_q^T(X\vert \vert Y),  \quad (q \neq 0,1 ),\\
&&D_q^R(X\vert \vert Y) = \frac{\log\left\{1+(q-1)D_q^T(X\vert \vert Y)\right\}}{q-1},\quad (q \neq 1), 
\end{eqnarray*}
where we set $q=\frac{1-\alpha}{2}$ in $D^{(q)}(X\vert \vert Y)$. Here, we note that
the form $\sum_{j=1}^n x_j^qy_j^{1-q}$ is appeared in all one-parameter extended relative entropies.
Therefore it was sufficient to study one quantity of them, for the study of a one-parameter extension of the relative entropy.
It is also notable that the Tsallis entropy can be rewritten by the Tsallis relative entropy as a special case:
$$
S_q(X) = \ln_q n -n^{1-q} D_{q}(X\vert\vert U)
$$
for the uniform distribution $U=\left\{1/n,\cdots,1/n\right\}$, where the $q$-logarithmic function is defined by 
$$
\ln_q(x) \equiv \frac{x^{1-q}-1}{1-q},\,\, q>0,q\ne 1,x>0.
$$
Thus the uniqueness theorem for the Tsallis relative entropy was proven in our previous paper \cite{Furu1}. 

In the present paper, as a further extension of our previous result, 
we give a two-parameter extended axiom for the function defined for any pairs of the probability distributions 
and prove the uniqueness theorem for a two-parameter extended relative entropy.
This paper is organized as follows.
In Section 2, we review the uniqueness theorem for relative entropy proven by A.Hobson, and 
the uniqueness theorem for a one-parameter extended relative entropy.
In Section 3, we show our main theorem. In Section 4, we characterize the constant appeared in Section 3.
In Section 5, we give properties for a two-parameter extended relative entropy.

\section{Review of the uniqueness theorem for one-parameter extended relative entropy}

The uniqueness theorem for relative entropy was shown by A. Hobson as follows \cite{Hob}:
\begin{The}{\bf (\cite{Hob})}\label{axiom_Hob}
The function $D_1(A\vert \vert B)$ is assumed to be defined for any two probability distributions $A=\left\{a_j\right\}$ and $B=\left\{b_j\right\}$ for $j=1,\cdots ,n$.
If $D_1(A \vert\vert  B)$ satisfies the following conditions (R1)-(R5), then it is given by the form $k \sum_{j=1}^na_j\log\frac{a_j}{b_j}$ with a positive constant $k$.
\begin{itemize}
\item[(R1)] {\it Continuity}: $D_1(A\vert \vert B)$ is a continuous function of $2n$ variables.
\item[(R2)] {\it Symmetry}: 
\begin{eqnarray*}
&&\hspace*{-18mm}  D_1\left(   a_1,\cdots ,a_j,\cdots ,a_k,\cdots ,a_n \vert \vert    b_1,\cdots ,b_j,\cdots ,b_k,\cdots ,b_n   \right) \\
&&\hspace*{-18mm}  =D_1\left(   a_1,\cdots ,a_k,\cdots ,a_j,\cdots ,a_n  \vert \vert    b_1,\cdots ,b_k,\cdots ,b_j,\cdots ,b_n   \right). 
\end{eqnarray*}
%\begin{eqnarray}
%&& D_1\left(   a_1,\cdots ,a_j,\cdots ,a_k,\cdots ,a_n \right. \nonumber \\
%&& \left.  \hspace*{8mm}  \vert \vert    b_1,\cdots ,b_j,\cdots ,b_k,\cdots ,b_n   \right) \nonumber \\
%&& = D_1\left(   a_1,\cdots ,a_k,\cdots ,a_j,\cdots ,a_n  \right. \nonumber \\ 
%&& \left. \hspace*{8mm}   \vert \vert    b_1,\cdots ,b_k,\cdots ,b_j,\cdots ,b_n   \right). 
%\end{eqnarray}
\item[(R3)] {\it Additivity}:
\begin{eqnarray*}
 &&  D_1 \left( a_{11} , \cdots ,a_{1m} ,a_{21} , \cdots ,a_{2m} \vert \vert  b_{11} , \cdots ,b_{1m} ,b_{21} , \cdots ,b_{2m}  \right)  = D_1 \left( {c_1 ,c_2 \left|\left|  {d_1 ,d_2 } \right.\right.    } \right) \\ 
 && + c_1 D_1 \left( {\frac{{a_{11} }}{{c_1 }}, \cdots ,\frac{{a_{1m} }}{{c_1 }} \left|\left|  {\frac{{b_{11} }}{{d_1 }}, \cdots ,\frac{{b_{1m} }}{{d_1 }}} \right.\right.  } \right)
+ c_2  D_1 \left( {\frac{{a_{21} }}{{c_2 }}, \cdots ,\frac{{a_{2m} }}{{c_2 }}\left|\left|  {\frac{{b_{21} }}{{d_2 }}, \cdots ,\frac{{b_{2m} }}{{d_2 }}} \right.\right.  } \right)  \label{Hob_H3}
 \end{eqnarray*}
where $c_i = \sum_{j=1}^m a_{ij}$ and $d_i = \sum_{j=1}^m b_{ij}$.
\item[(R4)]$D_1(A\vert \vert B) =0$ if $a_j =b_j$ for all $j$.
\item[(R5)]  $D_1(\frac{1}{n},\cdots ,\frac{1}{n},0,\cdots ,0\vert\vert  \frac{1}{n_0},\cdots ,\frac{1}{n_0})$ is an increasing function of $n_0$ and a decreasing function of $n$,
for any integers $n$ and $n_0$ such that $n_0 \geq n$.
\end{itemize}
\end{The}

As a one-parameter extension, we gave the uniqueness theorem for the Tsallis relative entropy as follows.
The function $D_q$ is defined for the probability distributions $A=\left\{a_j\right\}$ and $B=\left\{b_j\right\}$ on a finite probability space with one parameter $q \geq 0$. 
The one-parameter extended relative entropy (Tsallis relative entropy) was characterized by means of the following triplet of the generalized conditions (OR1), (OR2) and (OR3).  
\begin{Ax}{\bf (\cite{Furu1})} \label{g_hobson_axiom}

\begin{itemize}
\item[(OR1)] {\it Continuity}: $D_q(a_1,\cdots ,a_n \vert \vert  b_1, \cdots ,b_n)$ is a continuous function of $2n$ variables.
\item[(OR2)] {\it Symmetry}: 
\begin{eqnarray*}
&&\hspace*{-18mm}  D_q\left(   a_1,\cdots ,a_j,\cdots ,a_k,\cdots ,a_n \vert \vert    b_1,\cdots ,b_j,\cdots ,b_k,\cdots ,b_n   \right) \\
&&\hspace*{-18mm}  =D_q\left(   a_1,\cdots ,a_k,\cdots ,a_j,\cdots ,a_n  \vert \vert    b_1,\cdots ,b_k,\cdots ,b_j,\cdots ,b_n   \right). 
\end{eqnarray*}
%\begin{eqnarray*}
%&& D_q\left(   a_1,\cdots ,a_j,\cdots ,a_k,\cdots ,a_n   \right. \\
%&& \left. \hspace*{8mm}   \vert \vert    b_1, \cdots ,b_j,\cdots ,b_k,\cdots ,b_n   \right) \\
%&&= D_q\left(   a_1,\cdots ,a_k,\cdots ,a_j,\cdots ,a_n  \right. \\
%&& \left. \hspace*{8mm}   \vert \vert    b_1,  \cdots ,b_k,\cdots ,b_j,\cdots ,b_n   \right) 
%\end{eqnarray*}
\item[(OR3)] {\it Additivity}:
\begin{eqnarray}
&& D_q \left(   a_{11},\cdots ,a_{1m},\cdots ,a_{n1},\cdots ,a_{nm}  \vert \vert  b_{11},  \cdots ,b_{1m},\cdots ,b_{n1},\cdots ,b_{nm}   \right)   =  D_q(c_1,\cdots ,c_n \vert \vert  d_1 \cdots ,d_n)   \nonumber \\
&&      + \sum_{i=1}^n c_i^q d_i^{1-q} D_q\left( \frac{a_{i1}}{c_i},\dots ,\frac{a_{im}}{c_i} \left|\left|  \frac{b_{i1}}{d_i},\dots ,\frac{b_{im}}{d_i} \right.\right.  \right),\label{Tsa_H3}
\end{eqnarray} 
where $c_i = \sum_{j=1}^m a_{ij}$ and $d_i = \sum_{j=1}^m b_{ij}$.
\end{itemize}
\end{Ax}

Then, we have the following theorem.

\begin{The}\label{the-one-para} {\bf (\cite{Furu1})}
If conditions (OR1), (OR2) and (OR3) hold, then $D_q(A\vert B)$ is given in the following form:
\begin{equation}
D_q(A\vert\vert  B) = \sum_{j=1}^n  \frac{ a_j - a_j^q b_j^{1-q}}{\phi (q)}
\end{equation}
with a certain constant $\phi(q)$ depending on the parameter $q$. 
\end{The}

As for properties and applications of the Tsallis relative entropy, see our previous papers \cite{FYK,Furu2,Furu3}.

\section{Uniqueness theorem for two-parameter extended relative entropy} \label{sec3}
In our previous paper \cite{Furu1}, we gave Axiom \ref{g_hobson_axiom} in order to characterize the Tsallis relative entropy (one-parameter extended relative entropy).
In this section, we prove the uniqueness theorem for a two-parameter extended relative entropy.

\begin{The}\label{the1} 
If the function $D_{\alpha,\beta}(X\vert\vert  Y)$, defined for any pairs of the probability distributions $X=\left\{x_1,\cdots,x_n\right\}$ and $Y=\left\{y_1,\cdots,y_n\right\}$
on a finite probability space, satisfies the conditions (TR1)-(TR3) in the below, 
then  $D_{\alpha,\beta}(X \vert\vert  Y)$ is uniquely given by the form
\begin{equation}   \label{eq_the1}
D_{\alpha,\beta}(X\vert\vert  Y) = \sum_{j=1}^n  \frac{x_j^{\alpha}y_j^{1-\alpha} - x_j^{\beta} y_j^{1-\beta}}{\phi (\alpha,\beta)}
\end{equation}
with a certain constant $\phi(\alpha,\beta)$ depending on two parameters $\alpha$ and $\beta$.
\begin{itemize}
\item[(TR1)] {\it Continuity} : $D_{\alpha,\beta}  (x_1,\cdots ,x_n \vert \vert  y_1, \cdots ,y_n)$ is a continuous function for $2n$ variables.
\item[(TR2)] {\it Symmetry} : 
\begin{eqnarray*}
&&  D_{\alpha,\beta}(   x_1,\cdots ,x_j,\cdots ,x_k,\cdots ,x_n   \vert \vert   y_1,\cdots ,y_j,\cdots ,y_k,\cdots ,y_n   ) \\
&&    = D_{\alpha,\beta}(   x_1,\cdots ,x_k,\cdots ,x_j,\cdots ,x_n   \vert \vert  y_1,\cdots ,y_k,\cdots ,y_j,\cdots ,y_n   ). 
\end{eqnarray*}
\item[(TR3)] {\it Additivity} :
\begin{eqnarray}
&& D_{\alpha,\beta}(   x_{11},\cdots ,x_{1m},\cdots ,x_{n1},\cdots , x_{nm}   \vert \vert  y_{11},\cdots ,y_{1m},\cdots ,y_{n1},\cdots ,y_{nm}   ) \nonumber \\
&&  = D_{\alpha,\beta}(z_1,\cdots ,z_n \vert \vert  w_1 \cdots ,w_n) \sum_{j=1}^m \left(\frac{x_{ij}}{z_i}\right)^{\beta} \left(\frac{y_{ij}}{w_i}\right)^{1-\beta} \nonumber \\
&& + \sum_{i=1}^n z_i^{\alpha} w_i^{1-\alpha} D_{\alpha,\beta}\left( \frac{x_{i1}}{z_i},\dots ,\frac{x_{im}}{z_i} \left|\left|  \frac{y_{i1}}{w_i},\dots ,\frac{y_{im}}{w_i} \right.\right.  \right),\nonumber \\ \label{TR3}
\end{eqnarray} 
where $z_i = \sum_{j=1}^m x_{ij}$ and $w_i = \sum_{j=1}^m y_{ij}$.
\end{itemize}
\end{The}
{\it Proof}:
From (TR2), we have
\begin{eqnarray*}
 &&D_{\alpha ,\beta } \left( \frac{1}{{su}}, \cdots ,\frac{1}{{su}},0, \cdots ,0, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}},0, \cdots ,0 \left|\left|  \frac{1}{{tv}}, \cdots ,\frac{1}{{tv}} \right. \right.    \right) \\ 
 && = D_{\alpha ,\beta } \left( \frac{1}{{su}}, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}}, 0, \cdots ,0, \cdots ,0, \cdots ,0\left|\left|  \frac{1}{{tv}}, \cdots ,\frac{1}{{tv}} \right. \right.  \right). \\ 
 \end{eqnarray*}
From (TR3), we also have
\begin{eqnarray*}
&& D_{\alpha ,\beta } \left( \frac{1}{{su}}, \cdots ,\frac{1}{{su}},0, \cdots ,0, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}}, 0, \cdots ,0\left|\left|  \frac{1}{{tv}}, \cdots ,\frac{1}{{tv}} \right. \right.  \right) \\ 
&& \hspace*{-8mm}  = s\left( {\frac{1}{s}} \right)^\beta  \left( {\frac{1}{t}} \right)^{1 - \beta } D_{\alpha ,\beta } \left( {\frac{1}{u}, \cdots ,\frac{1}{u},0, \cdots ,0\left|\left|  {\frac{1}{v}, \cdots ,\frac{1}{v}} \right.\right.  } \right) \\
&& \hspace*{-8mm} + u\left( {\frac{1}{u}} \right)^\alpha  \left( {\frac{1}{v}} \right)^{1 - \alpha } D_{\alpha ,\beta } \left( {\frac{1}{s}, \cdots ,\frac{1}{s},0, \cdots ,0\left|\left|  {\frac{1}{t}, \cdots ,\frac{1}{t}} \right.\right.  } \right). \\ 
\end{eqnarray*}
From above two equations, we have
\begin{eqnarray*}
&&  D_{\alpha ,\beta } \left( \frac{1}{{su}}, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}}, \cdots ,\frac{1}{{su}}, 0, \cdots ,0, \cdots ,0, \cdots ,0\left|\left|  \frac{1}{{tv}}, \cdots ,\frac{1}{{tv}} \right. \right.  \right) \\ 
&&  = \left( {\frac{s}{t}} \right)^{1 - \beta } D_{\alpha ,\beta } \left( {\frac{1}{u}, \cdots ,\frac{1}{u},0, \cdots ,0\left|\left|  {\frac{1}{v}, \cdots ,\frac{1}{v}} \right.\right.  } \right) \\
&& + \left( {\frac{u}{v}} \right)^{1 - \alpha } D_{\alpha ,\beta } \left( {\frac{1}{s}, \cdots ,\frac{1}{s},0, \cdots ,0\left|\left|  {\frac{1}{t}, \cdots ,\frac{1}{t}} \right.\right.  } \right). \\ 
\end{eqnarray*}
If we put
\[
f_{\alpha ,\beta } \left( {s,t} \right) \equiv D_{\alpha ,\beta } \left( {\frac{1}{s}, \cdots ,\frac{1}{s},0, \cdots ,0 \left|\left|  {\frac{1}{t}, \cdots ,\frac{1}{t}} \right.  \right.} \right),\left( {t \ge s} \right),
\]
then we have
\[
f_{\alpha ,\beta } \left( {su,tv} \right) = \left( {\frac{s}{t}} \right)^{1 - \beta } f_{\alpha ,\beta } \left( {u,v} \right) + \left( {\frac{u}{v}} \right)^{1 - \alpha } f_{\alpha ,\beta } \left( {s,t} \right).
\]
We also have
\[
f_{\alpha ,\beta } \left( {us,vt} \right) = \left( {\frac{v}{u}} \right)^{1 - \beta } f_{\alpha ,\beta } \left( {s,t} \right) + \left( {\frac{s}{t}} \right)^{1 - \alpha } f_{\alpha ,\beta } \left( {u,v} \right),
\]
putting $s=u,u=s,t=v$ and $v=t$ in the above equation.
From above two equations, we have
\[
\frac{{\left( {\frac{s}{t}} \right)^{1 - \alpha }  - \left( {\frac{s}{t}} \right)^{1 - \beta } }}{{f_{\alpha ,\beta } \left( {s,t} \right)}} = \frac{{\left( {\frac{u}{v}} \right)^{1 - \alpha }  - \left( {\frac{u}{v}} \right)^{1 - \beta } }}{{f_{\alpha ,\beta } \left( {u,v} \right)}} \buildrel \Delta \over = \phi \left( {\alpha ,\beta } \right).
\]
Therefore we have
\[
f_{\alpha ,\beta } \left( {s,t} \right) = \frac{{\left( {\frac{s}{t}} \right)^{1 - \alpha }  - \left( {\frac{s}{t}} \right)^{1 - \beta } }}{{\phi \left( {\alpha ,\beta } \right)}}.
\]

For two natural numbers $l_i$ and $m_i$ such that $l_i \leq m_i$, we put
\[
z_i  \equiv \frac{{l_i }}{{\sum\limits_{k = 1}^n {l_k } }}, \left( {i = 1, \cdots ,n} \right),\quad 
w_i  \equiv \frac{{m_i }}{{\sum\limits_{k = 1}^n {m_k } }},   \left( {i = 1, \cdots ,n} \right)
\]
and
\begin{eqnarray*}
&&x_{ij}  \equiv \frac{1}{{\sum\limits_{k = 1}^n {l_k } }},\left( {i = 1, \cdots ,n;j = 1, \cdots ,l_i } \right),\\
&&y_{ij}  \equiv \frac{1}{{\sum\limits_{k = 1}^n {m_k } }},\left( {i = 1, \cdots ,n;j = 1, \cdots ,m_i } \right).
\end{eqnarray*}
From (TR2) and (TR3), we then have
\begin{eqnarray*}
&& D_{\alpha ,\beta } \left( \frac{1}{{\sum\limits_{k = 1}^n {l_k } }}, \cdots ,\frac{1}{{\sum\limits_{k = 1}^n {l_k } }},0, \cdots ,0, \cdots ,\frac{1}{{\sum\limits_{k = 1}^n {l_k } }}, \cdots ,\frac{1}{{\sum\limits_{k = 1}^n {l_k } }}, 0, \cdots ,0\left|\left|  \frac{1}{{\sum\limits_{k = 1}^n {m_k } }}, \cdots ,\frac{1}{{\sum\limits_{k = 1}^n {m_k } }} \right. \right. \right) \\ 
&& \hspace*{-8mm}   = D_{\alpha ,\beta } \left( {z_1 , \cdots ,z_n \left|\left|  {w_1 , \cdots ,w_n } \right.\right.} \right)\sum\limits_{j = 1}^{l_i } {\left( {\frac{1}{{l_i }}} \right)^\beta  \left( {\frac{1}{{m_i }}} \right)^{1 - \beta } } \\
&&\hspace*{-8mm}   + \sum\limits_{i = 1}^n {z_i^\alpha  w_i^{1 - \alpha } } D_{\alpha ,\beta } \left( {\frac{1}{{l_i }}, \cdots ,\frac{1}{{l_i }},0, \cdots ,0\left|\left|  {\frac{1}{{m_i }}, \cdots ,\frac{1}{{m_i }}} \right.\right. } \right), 
\end{eqnarray*}
since $x_{ij}=0$ for $j=l_i+1,\cdots ,m$.
Thus we have
\begin{eqnarray*}
&& D_{\alpha ,\beta } \left( {z_1 , \cdots ,z_n \left|\left|  {w_1 , \cdots ,w_n } \right.\right.} \right)  = \frac{{f_{\alpha ,\beta } \left( {\sum\limits_{k = 1}^n {l_k } ,\sum\limits_{k = 1}^n {m_k } } \right) - \sum\limits_{i = 1}^n {z_i^\alpha  w_i^{1 - \alpha } f_{\alpha ,\beta } \left( {l_i ,m_i } \right)} }}{{\sum\limits_{j = 1}^{l_i } {\left( {\frac{1}{{l_i }}} \right)^\beta  \left( {\frac{1}{{m_i }}} \right)^{1 - \beta } } }} \\ 
&& = \frac{  \left(\frac{  \sum_{k=1}^{n}l_k   }{\sum_{k=1}^{n}m_k   }\right)^{1-\alpha} -\left(\frac{  \sum_{k=1}^{n}l_k   }{\sum_{k=1}^{n}m_k   }\right)^{1-\beta}     }{\phi(\alpha,\beta)\left(\frac{l_i}{m_i}\right)^{1-\beta}} 
 - \frac{\sum_{i=1}^n z_i^{\alpha} w_i^{1-\alpha} \left\{ \left(\frac{l_i}{m_i}\right)^{1-\alpha} -\left(\frac{l_i}{m_i}\right)^{1-\beta}  \right\}}{\phi(\alpha,\beta)\left(\frac{l_i}{m_i}\right)^{1-\beta}} 
 \end{eqnarray*}
Here we have 
$$
\sum\limits_{i = 1}^n {z_i^r w_i^{1 - r} \left( {\frac{{l_i }}{{m_i }}} \right)^{1 - r} }   = \sum\limits_{i = 1}^n {\left( {\frac{{l_i }}{{\sum\limits_{k = 1}^n {l_k } }}} \right)^r \left( {\frac{{m_i }}{{\sum\limits_{k = 1}^n {m_k } }}} \right)^{1 - r} \left( {\frac{{l_i }}{{m_i }}} \right)^{1 - r} }  = \left( {\frac{{\sum\limits_{k = 1}^n {l_k } }}{{\sum\limits_{k = 1}^n {m_k } }}} \right)^{1 - r} ,\left( {r \in \mathbb{R}} \right)
$$
for 
\[
z_i  \equiv \frac{{l_i }}{{\sum\limits_{k = 1}^n {l_k } }},  \left( {i = 1, \cdots ,n} \right),\quad
w_i  \equiv \frac{{m_i }}{{\sum\limits_{k = 1}^n {m_k } }},  \left( {i = 1, \cdots ,n} \right).
\]
Thus we have
$$
D_{\alpha ,\beta } \left( {z_1 , \cdots ,z_n \left|\left|  {w_1 , \cdots ,w_n } \right.\right.    } \right) 
= \frac{{\sum\limits_{i = 1}^n {z_i^\alpha  w_i^{1 - \alpha } } \left( {\frac{{l_i }}{{m_i }}} \right)^{1 - \beta }  - \sum\limits_{i = 1}^n {z_i^\beta  w_i^{1 - \beta } \left( {\frac{{l_i }}{{m_i }}} \right)^{1 - \beta } } }}{{\phi \left( {\alpha ,\beta } \right)\left( {\frac{{l_i }}{{m_i }}} \right)^{1 - \beta } }}.
$$
Since we can take $l_i$ and $m_i$ arbitrary, we may take $l_i=l$ and $m_i=m$, then we have
\[
D_{\alpha ,\beta } \left( {z_1 , \cdots ,z_n \left|\left|  {w_1 , \cdots ,w_n } \right.\right.  } \right) 
= \frac{{\sum\limits_{i = 1}^n {z_i^\alpha  w_i^{1 - \alpha } }  - \sum\limits_{i = 1}^n {z_i^\beta  w_i^{1 - \beta } } }}{{\phi \left( {\alpha ,\beta } \right)}}.
\]
From (TR1) and the fact that any real number can be approximated by a rational number, the above result is true for any positive real number $z_j$ and $w_j$
satisfying $\sum_{j=1}^n z _j = \sum_{j=1}^n w_j =1$. 

\hfill \qed

Putting $\beta = 1$ and $\alpha=q$ in the above theorem, we have the uniqueness theorem for a one-parameter extended relative entropy (Theorem \ref{the-one-para}).

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Characterizations of $\phi(\alpha,\beta)$}
In this section, we characterize the constant $\phi(\alpha,\beta)$ depending on two parameters $\alpha$ and $\beta$.

\begin{Prop}
The postulate that our quantity $D_{\alpha ,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n } \right. \right.} \right)$ 
defined for any pairs of the probability distributions: 
\begin{equation} \label{postulate01}
D_{\alpha ,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.} \right)
 = \sum\limits_{j = 1}^n {\frac{{x_j^\alpha  y_j^{1 - \alpha }  - x_j^\beta  y_j^{1 - \beta } }}{{\phi \left( {\alpha ,\beta } \right)}}} 
\end{equation}
derived in Theorem \ref{the1} recovers the relative entropy
 when $\alpha \to 1$ and $\beta \to 1$, that is,
\begin{equation} \label{postulate02}
\mathop {\lim }\limits_{\alpha ,\beta  \to 1} D_{\alpha ,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.} \right)
 = k\sum\limits_{j = 1}^n {x_j \left(\log x_j -\log y_j\right)} 
\end{equation}
implies the following conditions.
\begin{itemize}
\item[(c1)] We have $
\mathop {\lim }\limits_{\alpha  \to 1} \phi \left( {\alpha ,1} \right) = \mathop {\lim }\limits_{\beta  \to 1} \phi \left( {1,\beta } \right) 
= \mathop {\lim }\limits_{\beta  \to \alpha } \phi \left( {\alpha ,\beta } \right) = 0$
and $\phi \left( {\alpha ,\beta } \right) \ne 0\,\,for\,\,\alpha  \ne \beta$.
\item[(c2)] There exists the interval $(a,b)$ such that $\phi(\alpha,1)$  and  $\phi(1,\beta)$ are differentiable on  $\left( {a,1} \right) \cup \left( {1,b} \right)$.
\item[(c3)]There exists the constant $k>0$  such that  $
\mathop {\lim }\limits_{\alpha  \to 1} \frac{{d\phi \left( {\alpha ,1} \right)}}{{d\alpha }} = \frac{1}{k}$
and
$
\mathop {\lim }\limits_{\beta  \to 1} \frac{{d\phi \left( {1,\beta } \right)}}{{d\beta }} =  - \frac{1}{k}$.
\end{itemize}
\end{Prop}

{\it Proof}:
\begin{itemize}
\item[(c1)] 
We may calculate the limit of the left hand side in Eq.(\ref{postulate02}) in the following ways.
\begin{itemize}
\item[(i)]
Firstly we may take the limit $\alpha\to 1$ in Eq.(\ref{postulate01}) 
and then later take the limit $\beta \to 1$:
$$
\hspace*{-15mm} \mathop {\lim }\limits_{\beta  \to 1} D_{1,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.} \right)
 = \mathop {\lim }\limits_{\beta  \to 1} \sum\limits_{j = 1}^n {\frac{{x_j  - x_j^\beta  y_j^{1 - \beta } }}{{\phi \left( {1,\beta } \right)}}}. 
$$
Since we have
$\mathop {\lim }\limits_{\beta  \to 1} \sum\limits_{j = 1}^n {\left( {x_j  - x_j^\beta  y_j^{1 - \beta } } \right)}  = 0$, 
we need $\mathop {\lim }\limits_{\beta  \to 1} \phi \left( {1,\beta } \right) = 0$ in order that we have the limit in the above.
\item[(ii)] By the similar way to (i), we have $\mathop {\lim }\limits_{\alpha  \to 1} \phi \left( {\alpha ,1} \right) = 0$.
\item[(iii)] Firstly we may put $\beta \to \alpha$ 
and then later take the limit $\alpha\to 1$.
In the case $\beta \to \alpha$, the summation of the numerator of the right hand side in Eq.(\ref{postulate01}) is equal to $0$:
$$\lim_{\beta \to \alpha}\sum_{j=1}^n \left(x_j^{\alpha} y_j^{1-\alpha}-x_j^{\beta} y_j^{1-\beta}\right)=0.$$ 
Therefore we have $\lim_{\beta\to\alpha}\phi(\alpha,\beta)=0$, otherwise 
$\lim_{\beta\to\alpha}D_{\alpha,\beta}(x_1,\cdots,x_n\vert\vert  y_1,\cdots,y_n)$ takes $0$, which contradicts the Eq.(\ref{postulate02}).
From the reason why we have the limit of the left hand side in (\ref{postulate02}),
 we also have $\phi \left( {\alpha ,\beta } \right) \ne 0$ for $ \alpha \neq \beta$, since $
\sum\limits_{j = 1}^n {x_j^\alpha  y_j^{1 - \alpha }  - x_j^\beta  y_j^{1 - \beta } }  \ne 0$ for $ \alpha \neq \beta$.
\end{itemize}
\item[(c2)] Since $
\sum\limits_{j = 1}^n {x_j  - x_j^\beta  y_j^{1 - \beta } } $ is differentiable by $\beta$, we need that there exists an interval $(a,b)$ such that
$\phi(1,\beta)$ is also differentiable by $\beta$ on $\left( {a,1} \right) \cup \left( {1,b} \right)$, in order that we have the limit of the left hand side in Eq.(\ref{postulate02}).
By the similar way,  there exists an interval $(a,b)$ such that
$\phi(\alpha,1)$ is also differentiable by $\beta$ on $\left( {a,1} \right) \cup \left( {1,b} \right)$.
\item[(c3)] Since we have
\begin{eqnarray*}
\mathop {\lim }\limits_{\beta  \to 1} D_{1,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.  } \right) 
&=& \mathop {\lim }\limits_{\beta  \to 1} \sum\limits_{j = 1}^n {\frac{{x_j  - x_j^\beta  y_j^{1 - \beta } }}{{\phi \left( {1,\beta } \right)}}}\\
&=& \mathop {\lim }\limits_{\beta  \to 1} \frac{{ - \sum\limits_{j = 1}^n {x_j^\beta  y_j^{1 - \beta } \left( {\log x_j  - \log y_j } \right)} }}{{\frac{{d\phi \left( {1,\beta } \right)}}{{d\beta }}}},
\end{eqnarray*} 
there exists a constant $k>0$ such that $
\frac{{d\phi \left( {1,\beta } \right)}}{{d\beta }} =  - \frac{1}{k}$.
By the similar way, 
there exists a constant $k>0$ such that $
\frac{{d\phi \left( {\alpha,1 } \right)}}{{d\beta }} =   \frac{1}{k}$.
\end{itemize}

\hfill \qed

\begin{Prop}
$D_{\alpha,\beta}(X \vert\vert  U)$ takes the minimum value for fixed posterior probability distribution as uniform distribution 
$U=\left\{\frac{1}{n},\cdots ,\frac{1}{n}  \right\}$ :
$$D_{\alpha,\beta}\left(x_1,\cdots ,x_n \left|\left|  \frac{1}{n},\cdots ,\frac{1}{n}\right.\right.  \right) 
\geq D_{\alpha,\beta}\left(\frac{1}{n},\cdots ,\frac{1}{n}  \left|\left|  \frac{1}{n},\cdots ,\frac{1}{n}\right.\right.  \right), $$
when we have  
\begin{itemize}
\item[(c4)] the following relations  (i) and (ii) for $\alpha$ and $\beta$
\begin{itemize}
\item[(i)] $\alpha\neq\beta$.
\item[(ii)]
  If $\phi \left( {\alpha ,\beta } \right) > 0$, then we have $0 \le \beta  \le 1 \le \alpha$.  
  If $\phi \left( {\alpha ,\beta } \right) < 0$, then we have $0 \le \alpha  \le 1 \le \beta$. 
\end{itemize}

\end{itemize}

\end{Prop}
{\it Proof}:
The second derivative of $
D_{\alpha ,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {\frac{1}{n}, \cdots ,\frac{1}{n}} \right.\right.} \right)
$ on $x_j$ is calculated by 
$$
 \frac{{d^2 D_{\alpha ,\beta } \left( {x_1 , \cdots ,x_n \left|\left|  {\frac{1}{n}, \cdots ,\frac{1}{n}} \right.\right.} \right)}}{{dx_j^2 }}  = \frac{{n^{\alpha  - 1} \alpha \left( {\alpha  - 1} \right)x_j^{\alpha  - 2}  - n^{\beta  - 1} \beta \left( {\beta  - 1} \right)x_j^{\beta  - 2} }}{{\phi \left( {\alpha ,\beta } \right)}}
$$
This takes positive value in the case of (c4) so that it should be convex in $x_j$. Therefore $D_{\alpha,\beta}(X\vert\vert  U)$ takes the minimum value. 

\hfill \qed

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Properties of a two-parameter extended relative entropy}
As an example satisfying  the conditions (c1)-(c4) on $\phi(\alpha,\beta)$, we simply take $\phi(\alpha,\beta) = \alpha -\beta$.
Then we may define a two-parameter extended relative entropy in the following.
\begin{Def}
For two parameters $\alpha,\beta\in  \mathbb{R}$ satisfying $ 0 \le \alpha  \le 1 \le \beta $
or  $0 \le \beta  \le 1 \le \alpha$, and 
two probability distributions $X=\left\{x_1,\cdots,x_n\right\}$ and $Y=\left\{y_1,\cdots,y_n\right\}$,
we define a two-parameter extended relative entropy by
\[
D_{\alpha ,\beta } \left( {X \left| \left|  {Y} \right.\right.} \right) 
\equiv \sum\limits_{j = 1}^n {\frac{{x_j^\alpha  y_j^{1 - \alpha }  - x_j^\beta  y_j^{1 - \beta } }}{{\alpha  - \beta }}},\quad (\alpha\neq\beta). 
\]
\end{Def}
Note that a two-parameter extended relative entropy is a generalization of the relative entropy in the sense that
$$
\lim_{\alpha,\beta\to 1}D_{\alpha ,\beta } \left( {X \left|\left|  {Y} \right.\right.} \right) = D_1(X\vert \vert Y).
$$
We also note that a two-parameter extended relative entropy recovers 
the Tsallis relative entropy (one-parameter extended relative entropy)
when $\alpha=1$ or $\beta=1$. 
The Tsallis relative entropy is also a one-parameter generalization of the relative entropy:
$$\lim_{q\to 1}D_{q}^T(X \vert \vert  Y) = D_1(X\vert \vert Y).$$

In addition, we note that a two-parameter extended relative entropy  is expressed by the convex combination of the Tsallis relative entropy:
\begin{equation}\label{relation}
D_{\alpha,\beta}(X \vert \vert  Y) = \frac{\alpha-1}{\alpha-\beta}D_{\alpha}^T(X \vert \vert  Y)+\frac{1-\beta}{\alpha-\beta}D_{\beta}^T(X \vert \vert  Y).
\end{equation}

Thus we have the following properties on a two-parameter relative entropy, thanks to the above relation and the properties of the Tsallis relative entropy studied in \cite{FYK}.
\begin{Prop}\label{prop_1}
For a two-parameter extended relative entropy $D_{\alpha,\beta}(X \vert \vert  Y)$, we have the following properties.
\begin{itemize}
\item[(i)] (Nonnegativity) $D_{\alpha,\beta}  (X \vert \vert  Y) \geq 0$.
\item[(ii)] (Symmetry) 
$$
 D_{\alpha,\beta} \left( {x_{\pi \left( 1 \right)} , \cdots ,x_{\pi \left( n \right)} \left|\left|  {y_{\pi \left( 1 \right)} , \cdots ,y_{\pi \left( n \right)} } \right.\right.} \right) 
 = D_{\alpha,\beta} \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.} \right).
$$
\item[(iii)] (Possibility of extension) 
$$
 D_{\alpha,\beta} \left( {x_1 , \cdots ,x_n ,0 \left|\left|  {y_1 , \cdots ,y_n,0} \right.\right.} \right) 
 = D_{\alpha,\beta} \left( {x_1 , \cdots ,x_n \left|\left|  {y_1 , \cdots ,y_n} \right.\right.} \right).
$$
\item[(iv)] (Joint convexity) For $0 \leq \lambda \leq 1$ and the probability distributions $X^{(i)}=\left\{x_j^{(i)}\right\}$,$Y^{(i)}=\left\{y_j^{(i)}\right\}$, $(i=1,2;j=1,\cdots,n)$,  we have 
\begin{eqnarray*}
\hspace*{-10mm}  &&D_{\alpha,\beta} \left( \lambda X^{\left( 1 \right)}  + \left( {1 - \lambda } \right)X^{\left( 2 \right)} \left|\left|  \lambda Y^{\left( 1 \right)}  + \left( {1 - \lambda } \right)Y^{\left( 2 \right)}  \right.\right.\right) \\
\hspace*{-10mm} &&\le \lambda D_{\alpha,\beta} \left( {X^{\left( 1 \right)} \left|\left|  Y^{\left( 1 \right)} \right.\right. } \right) + \left( {1 - \lambda } \right)D_{\alpha,\beta} \left( {X^{\left( 2 \right)} \left|\left|  Y^{\left( 2 \right)} \right.\right.} \right).
\end{eqnarray*}
\item[(v)] (Monotonicity)  For the transition probability matrix $W$, we have
\[D_{\alpha,\beta} \left( {WX \left| \left| {WY} \right.\right.} \right) \le D_{\alpha,\beta} \left( {X \left|\left|  Y \right.\right.} \right).\]
\end{itemize}
\end{Prop}

It is also notable that we have the following expression for a two-parameter extended relative entropy:
$$
D_{\alpha,\beta}(X\vert\vert U) = \frac{n^{\alpha-1}-n^{\beta -1}}{\alpha-\beta} -n^{\alpha -1} \left( \frac{\alpha -1}{\alpha-\beta}\right) S_{\alpha}(X) -
n^{\beta -1} \left(\frac{1-\beta}{\alpha-\beta} \right)S_{\beta}(X) 
$$
for the uniform distribution $U=\left\{1/n,\cdots,1/n\right\}$,
while we also have the following relation between the two-parameter extended entropy and the Tsallis entropy (one-parameter extended entropy):
$$
S_{\alpha,\beta}(X) = \left( \frac{\alpha -1}{\alpha-\beta}\right) S_{\alpha}(X) + \left( \frac{1-\beta}{\alpha-\beta} \right) S_{\beta}(X). 
$$
Therefore we may not obtain the direct relation between $D_{\alpha,\beta}(X\vert\vert U)$ and $S_{\alpha,\beta}(X)$ except for $n=1$.
%However, if we consider a simple distribution such as $I=\left\{ 1,\cdots,0\right\}$ for example, we easily find their relation as
%$$S_{\alpha,\beta}(X) = - D_{\alpha,\beta}(X\vert\vert I). $$

\section{Conclusion} 

As we have seen in Section \ref{sec3}, the two-parameter extended relative entropy is characterized by continuity, symmetry and additivity.
On the other hand, it is known that the $f$-divergence is characterized by symmetry, monotonicity and joint convexity \cite{Csi2}. 
The properties such as monotonicity and joint convexity are represented by the inequalities.
For the characterization of $f$-divergence, we need the inequalities involving their equality conditions, while for the characterization
of  the two-parameter extended relative entropy, we need the functional equation referred by an additivity.  
Therefore the conditions in our axiom are essentially different from those of the axiom characterizing $f$-divergence. 
It is also notable that our characterization of a two-parameter extended relative entropy (Theorem 3.1) is the uniqueness theorem
such that the function $D_{\alpha,\beta}(X\vert\vert Y)$ is uniquely given by Eq.(\ref{eq_the1}), while
the characterization of $f$-divergence (Theorem 1 in \cite{Csi2}) is the existence theorem for a convex function $f$ such that
the function defined for any pair of the probability distributions is equal to the $f$-divergence.
In other words, in the paper \cite{Csi2}, the  existence of the convex function has been shown but
 the uniqueness of the convex function $f$ has not been shown, 
so that our axiomatic characterization may have an advantage since it uniquely gives a two-parameter extended relative entropy.
It is also notable that the uniqueness theorem for $\alpha$-divergence was recently shown in \cite{Ama} for the special case such that
the divergence measure (functional) is written by a sum of all components.

Closing this section, we give the expressions of a two-parameter extended relative entropy by means of $f$-divergence:
$$
D_f(X\vert\vert Y)\equiv \sum_{j=1}^n y_jf\left(\frac{x_j}{y_j}\right),
$$
where $f$ is a convex function on $(0,\infty)$ and $f(1)=0$.
If we take $f(t)=t \log t$, then $f$-divergence $D_f(X\vert\vert Y)$ recovers the relative entropy.
Here, if we put 
\begin{equation} \label{f-div_func01}
f_{\alpha,\beta}(t) \equiv \frac{t^{\alpha}-t^{\beta}}{\alpha-\beta}, \quad (\alpha \neq \beta),
\end{equation}
then
$\frac{d^2f_{\alpha,\beta}(t)}{dt^2} \geq 0$ for $0 \le \alpha  \le 1 \le \beta $
or  $0 \le \beta  \le 1 \le \alpha$.
And then we have the following expression:
$$
D_{\alpha,\beta}(X\vert\vert Y) = D_{f_{\alpha,\beta}}(X\vert\vert Y).
$$

It is known that the relative entropy is connected to many important results in the mathematical physics and information science.
For a two-parameter extended relative entropy, such connections (for example with H-theorem or variational expressions related to the free energy) will be studied in the future.

%The $f$-divergence is often defined by
%$$D_{f^*}(X\vert\vert Y)\equiv \sum_{j=1}^n x_jf^*\left(\frac{y_j}{x_j}\right),$$
%where $f$ is a convex function on $(0,\infty)$ and $f(1)=0$.
%If we take $f^*(t)=- \log t$, then $f$-divergence $D_{f^*}(X\vert\vert Y)$ recovers the relative entropy.
%Here, if we put 
%\begin{equation} \label{f-div_func02}
%f^{*}_{\alpha,\beta}(t) \equiv \frac{t^{1-\alpha}-t^{1-\beta}}{\alpha-\beta}, \quad (\alpha \neq \beta),
%\end{equation}
%then $\frac{d^2f^{*}_{\alpha,\beta}(t)}{dt^2} \geq 0$ for $0 \le \alpha  \le 1 \le \beta $or  $0 \le \beta  \le 1 \le \alpha$.
%Indeed, if the function  $f(t)$ is convex on $(0,\infty)$, then the function $g(t)\equiv t f\left(\frac{1}{t}\right)$ is also convex on $(0,\infty)$, because of the elementary calculation:
%$$\frac{d^2g(t)}{dt^2} = \frac{1}{t^3}\frac{d^2f\left(\frac{1}{t}\right)}{dt^2}.$$
%And then we have the following expression:
%$$D_{\alpha,\beta}(X\vert\vert Y) = D_{f^{*}_{\alpha,\beta}}(X\vert\vert Y).$$
%It is known that the dual of $f$-divergence:
%$$D^*_f(X\vert\vert Y) \equiv D_f(Y\vert \vert X)$$
%has the relation 
%$$D^*_f(X\vert\vert Y) = D_{f^*}(X\vert \vert Y),$$
%if we have $f^*(t)=tf(\frac{1}{t})$. (For example, see \cite{AN} for details.)
%In the case of two-parameter extended relative entropy, the above relation holds:
%$f^*_{\alpha,\beta}(t) = t f_{\alpha,\beta} (\frac{1}{t})  $.
%Through the concept of duality in the field of information geometry, we find the relation between Eq.(\ref{f-div_func01}) and Eq.(\ref{f-div_func02}).

\section*{Acknowledgements}
I  would like to thank Professor H.Suyari and Professor T.Wada giving me an opportunity to read their interesting paper \cite{WS} in the workshop at Chiba University.
The author was partially supported by the Japanese Ministry of Education, Science, Sports and Culture, 
Grant-in-Aid for Encouragement of Young Scientists (B) 20740067.

\end{document}